LUCIFER Commissioning (1a) and AGw Commissioning for 17-18 September 2008 (Day 19)

Observers: N. Ageorges, W. Seifert, and numerous other LUCIFERians (Peter, Kai, Marcus, Volker, Andreas) on mtn, D. Thompson.
J.Hill and D. Miller in Tucson
Software Support: P. Grenz, N. Cushing, C. Biddick, M. De La Pena in Tucson
Telescope Operators: A. Ceranski

Summary

Good weather and no software or hardware problems after the first 2 hours of the night allowed Lucifer to get a good solid night of commissioning data. People on the mountain and in Tucson were quite happy!

The first 1.5 hours were spent installing a new version of the IRS, and realizing the GCS GUI was causing GCS to crash. Once these problems were found and fixed, the GCS only died once during the night and the IRS never locked up.

The Guider software worked quite well, with only (I think) one failure during the evening. This was due to the fact that we had not reset the pointing zero points for several hours. The fix took about 7-8 minutes and did not have to be reset for the rest of the night (~3 hours).

The Active Optics loop worked very well. It always started, ran, paused, restarted as requested.

The Active Optics corrections where very good 75% of the time, on and off-axis. There were several instances were the Lucifer images were showing aberration (sausages and some small tails) A couple of these times were when we were at low (40) elevation, 10 second WFSC integrations (bright guide star!), pupil offset not corrected properly (will be fixed by tonight) and 4.5' off-axis guide star. All these issues need to be addressed, but the near term solution is to choose field and guide stars so these problem areas are avoided. A couple times it is not clear why the corrections were not good, but may be caused by telescope vibrations. I would estimate, if it were vibrations, that it occurred for about 30 - 40 minutes during the night.

At the end of the night, it was discovered that the first Lucifer image in a sequence after a OffsetTelscope request had stars with either trails (up to 4" long) or looked like sausages (sometimes with a ratio of 3-to-1). It was at first thought that Lucifer was taking images to soon and the telescope was still in motion, but test data with various wait times suggest that this is not the problem. More investigation tonight.

Nighttime Activities

2:15 Opened at sunset

SW group worked on IRS. Installed new version of IRS.

4:00 PresetTelescope to Lucifer star, on-axis. Collimate wfsc 1-18

4:20 PresetTelescope to off-axis star on same field about 3'.
Collimate wfsc 19-35-47
Send a couple hundred nm of focus on first iteration, but after that usually less than 100 nm corrections

4:46 PresetTelescope to new field, off-axis, 2'
First collimation, 48 sent focus = 900, coma = 500
Next iterations about 100 nm
wfsc 48-84

5:17 Preset to new field. GCS died. Aaron restarted
. Re-send preset. Guide correction seem to not be applied on the first image. Star was along the bottom edge so may have calculated a 0 correction for one iteration.
wfsc off-axis ~1.5" wfsc 85-93

5:20 A LUCIFERian passed out. A little O2 and is feeling better

5:23 Aaron move telescope to Zenith and went to SX upper treehouse, left of door, labeled Laser-Tracker, rotary switch

5:35 Preset to same field. After several minutes and good images, a 10" offset was sent and images, 4 over 10 seconds, appeared to have some astigmatism. Collimate looks good from wfsc view. Vibrations?

5:45 New preset. wfsc ~111-~160

6:40 Chris and Norm find that the isGuiding and isWFSing DD flags are not reset back to true after an offset request. IssueTrak #1646

===> Lucifer happily continues taking data

12:30 Lucifer done taking data and the telescope was closed.

-- DougMiller - 18 Sep 2008
Topic revision: r4 - 18 Sep 2008, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback